# Cosmos World Foundation Model Platform for Physical AI

### v1

@@ 5.2.6 Results
In Fig. 17, we show qualitative results of our autoregressive WFMs using different model sizes. In the unprompted setting, comparing Cosmos-1.0-Autoregressive-4B and Cosmos-1.0-Autoregressive-12B model, we observe that the 12B model generates videos with better motion and sharper details. Similarly, in the prompted setting, comparing Cosmos-1.0-Autoregressive-5B-Video2World and Cosmos-1.0-Autoregressive-13B-Video2World reveals that the 13B model gets better motion than the 5B model.
@@
| Output of Cosmos-1.0-Autoregressive-4B | |||
@@
| Cosmos-1.0-Autoregressive-4B | ||
| Cosmos-1.0-Autoregressive-5B-Video2World | ||
| Cosmos-1.0-Autoregressive-12B | ||
| Cosmos-1.0-Autoregressive-13B-Video2World |
@@ 5.2.7 Limitations
One notable failure case observed in the generated videos of our autoregressive WFMs is objects unexpectedly appearing from below. Fig. 19 illustrates an example of this issue. To understand the failure rate of our models, we conduct a systematic study by creating an evaluation set of Physical AI inputs to our autoregressive WFMs. We generate videos with all our models using two input modes—image (single-frame) conditioning and video (9-frame) conditioning. For all generated videos, we manually inspect the failure cases and report the failure rate in Tab. 18. We observe that the smaller models Cosmos-1.0-Autoregressive-4B and Cosmos-1.0-Autoregressive-5B-Video2World show a higher corruption rate in single frame conditioning, while the larger models Cosmos-1.0-Autoregressive-12B and Cosmos-1.0-Autoregressive-13B-Video2World are more robust. Generation with -frame video conditioning is stable for all models, with a failure rate lower than .
@@ 5.3 Evaluation
| Cosmos-1.0-Diffusion-7B-Text2World | 0.355 | 62.6% | 33.02 | 0.939 | 0.070 |
| Cosmos-1.0-Diffusion-7B-Video2World | 0.473 | 68.4% | 30.66 | 0.929 | 0.085 |
| Cosmos-1.0-Autoregressive-4B | 0.433 | 35.6% | 32.56 | 0.933 | 0.090 |
| Cosmos-1.0-Autoregressive-5B-Video2World | 0.392 | 27.0% | 32.18 | 0.931 | 0.090 |
@@ 5.3.2 Physics Alignment
| Cosmos-1.0-Diffusion-7B-Video2World | prompt + 1 frame | 17.34 | 0.538 | 0.836 | 0.332 |
| Cosmos-1.0-Diffusion-7B-Video2World | prompt + 9 frames | 21.06 | 0.691 | 0.859 | 0.592 |
| Cosmos-1.0-Diffusion-14B-Video2World | prompt + 1 frame | 16.81 | 0.521 | 0.836 | 0.338 |
| Cosmos-1.0-Diffusion-14B-Video2World | prompt + 9 frames | 20.21 | 0.635 | 0.860 | 0.598 |
| Cosmos-1.0-Autoregressive-4B | 1 frame | 17.91 | 0.486 | 0.827 | 0.394 |
| Cosmos-1.0-Autoregressive-4B | 9 frames | 18.13 | 0.482 | 0.859 | 0.481 |
| Cosmos-1.0-Autoregressive-5B-Video2World | prompt + 1 frame | 17.67 | 0.478 | 0.818 | 0.376 |
| Cosmos-1.0-Autoregressive-5B-Video2World | prompt + 9 frames | 18.29 | 0.481 | 0.864 | 0.481 |
| Cosmos-1.0-Autoregressive-12B | 1 frame | 17.94 | 0.486 | 0.829 | 0.395 |
| Cosmos-1.0-Autoregressive-12B | 9 frames | 18.22 | 0.487 | 0.869 | 0.487 |
| Cosmos-1.0-Autoregressive-13B-Video2World | prompt + 1 frame | 18.00 | 0.486 | 0.830 | 0.397 |
| Cosmos-1.0-Autoregressive-13B-Video2World | prompt + 9 frames | 18.26 | 0.482 | 0.865 | 0.482 |
@@ 6 Post-trained World Foundation Model
| Sec. 6.1 | Cosmos-1.0-Diffusion-7B-Video2World-Sample-CameraCond | Text + Image + Cameras |
| Sec. 6.2 | Cosmos-1.0-Autoregressive-7B-Video2World-Sample-Instruction | Text + Video |
| Sec. 6.2 | Cosmos-1.0-Diffusion-7B-Video2World-Sample-Instruction | Text + Video |
| Sec. 6.2 | Cosmos-1.0-Autoregressive-7B-Video2World-Sample-ActionCond | Action + Video |
| Sec. 6.2 | Cosmos-1.0-Diffusion-7B-Video2World-Sample-ActionCond | Action + Video |
| Sec. 6.3 | Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView | Text |
| Sec. 6.3 | Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond | Text + Trajectory |
| Sec. 6.3 | Cosmos-1.0-Diffusion-7B-Video2World-Sample-MultiView | Text + Video |
@@ 6.1 Post-training WFM for Camera Control
Through camera pose conditioning, we integrate camera control into Cosmos-1.0-Diffusion-7B-Video2World, making it an effective 3D world simulator. We term the result post-trained WFM as Cosmos-1.0-Diffusion-7B-Video2World-Sample-CameraCond. We focus on generating 3D worlds from a single reference input image, leveraging camera control to produce temporally coherent and 3D-consistent video simulations from the specified camera trajectories, where changes in perspective align with the underlying 3D structure of the scene.
@@ 6.1.2 Fine-tuning
where is the camera center location and is the unit ray direction of each latent pixel (where the latent embedding is treated as a downsampled image). All the camera poses are relative with respect to the initial frame. The Cosmos-1.0-Tokenizer-CV8x8x8 used by Cosmos-1.0-Diffusion-7B-Video2World models has a temporal compression rate of , and thus for every 8 frames, we use the Plücker embedding at the 4th frame to concatenate with the corresponding latent representation.
@@ 6.1.3 Evaluation
| Cosmos-1.0-Diffusion-7B-Video2World- | 82.0% | 1.646 | 0.038 | 14.30 | 120.49 |
@@
Comparisons. We present the results in Tab. 22. First, our post-trained WFM can generate realistic and coherent 3D worlds. This is evidenced by the lower FID/FVD scores (higher visual quality) and the higher camera pose estimation success rate. Cosmos-1.0-Diffusion-7B-Video2World-Sample-CameraCond demonstrates better camera control, as the camera trajectory re-estimation is significantly closer to the original control input.
We also provide visual comparisons in Fig. 21. While CamCo struggles to generate content beyond the input image, Cosmos-1.0-Diffusion-7B-Video2World-Sample-CameraCond effectively generates visuals that adhere to the structure of a 3D world. Note that both models were post-trained on DL3DV-10K and evaluated on the RealEstate10K dataset, which introduces a significant distribution shift between training and testing. The Cosmos model successfully overcomes this distribution shift while also demonstrating its capability to generalize to unseen input camera trajectories.
@@
To show the diversity of the generation, we show generation results from the same input image and camera control with different random seeds in Fig. 23. Cosmos-1.0-Diffusion-7B-Video2World-Sample-CameraCond is able to generate different worlds while still maintaining 3D spatial and temporal coherence in the videos. This could be used to simulate different possible futures given the current states.
@@ 6.2.2 Fine-tuning
We fine-tune both our Cosmos-1.0-Diffusion-7B-Video2World (Sec. 5.1) and Cosmos-1.0-Autoregressive-5B-Video2World (Sec. 5.2) for instruction-based video prediction and action-based next-frame prediction tasks.
For instruction-based video prediction, we build two models based on the base WFMs. The first is called Cosmos-1.0-Diffusion-7B-Video2World-Sample-Instruction, and the second is called Cosmos-1.0-Autoregressive-5B-Video2World-Sample-Instruction. We compute the T5 embedding of the instruction, which is added to the finetuing of the base model via cross-attention.
For action-based next-frame prediction, we also build two models based on the base WFMs. The first one is called Cosmos-1.0-Diffusion-7B-Video2World-Sample-ActionCond, and the second one is called Cosmos-1.0-Autoregressive-5B-Video2World-Sample-ActionCond.
Since action is a new modality not encountered during pre-training, we introduce additional modules inside our models for conditioning. For Cosmos-1.0-Autoregress-5B-Video2World-Sample-ActionCond, we add an action embedder MLP to project the action vector into a tensor, which is then incorporated into the model via cross-attention. For Cosmos-1.0-Diffusion-7B-Video2World-Sample-ActionCond, we also add an action embedder MLP to predict the action into a tensor but instead, incorporate it into the model by adding it to the timestamp embedding of the DiT modules.
@@ 6.2.3 Evaluation
| Cosmos-1.0-Diffusion-7B-Video2World-Sample-Instruction | Cosmos-1.0-Autoregressive-5B-Video2World-Sample-Instruction | |||||||||
@@
| Cosmos-1.0-Diffusion-7B-Video2World-Sample-ActionCond | Cosmos-1.0-Autoregressive-5B-Video2World-Sample-ActionCond | ||||||||||
@@
As shown, we find that both Cosmos-1.0-Diffusion-7B-Video2World-Sample-Instruction and Cosmos-1.0-Autoregressive-5B-Video2World-Sample-Instruction perform better than VideoLDM-Instruction along the four evaluation dimensions. Cosmos-1.0-Diffusion-7B-Video2World-Sample-Instruction achieved overall preference compared to for VideoLDM-Instruction. Cosmos-1.0-Autoregressive-5B-Video2World-Sample-Instruction has also achieved better performance than diffusion-based VideoLDM-Instruction. Some predicted video frames for both fine-tuned WFMs are presented in Fig. 25, which shows the quality of the predicted videos.
@@
| Cosmos-1.0-Autoregressive-5B-Video2World- Sample-ActionCond | 19.95 | 0.80 | 0.36 | 434 |
| Cosmos-1.0-Diffusion-7B-Video2World- Sample-ActionCond | 21.14 | 0.82 | 0.32 | 190 |
The computed metrics are summarized in Tab. 23, including PSNR, SSIM, Latent L2 (Zhu et al., 2024), and FVD. As shown, both Cosmos-1.0-Autoregressive-5B-Video2World-Sample-ActionCond and Cosmos-1.0-Diffusion-7B-Video2World-Sample-ActionCond models outperform the baseline model (IRASim-Action). Some predicted video frames are presented in Fig. 26, which shows the quality of the predicted videos compared to the ground truth.
@@ 6.3.2 Fine-tuning
We fine-tune our Cosmos-1.0-Diffusion-7B-Text2World (Sec. 5.1) into a multiple-view world model using the RDS dataset. To ensure consistent video generation across multiple views, we slightly modify the architectural design described in Sec. 5.1 and fine-tune the WFM to generate videos from all six cameras simultaneously.
We build three multi-view world models, summarized in Tab. 21. The first one is called Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView, which is a multi-view world model that can generate six camera views based on a text prompt input. The second one is called Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond. This model is built on top of Cosmos-1.0-Diffusion-7B-Text2World-Sample-AV-MultiView and takes an additional trajectory input as the conditional input signal. The final model, Cosmos-1.0-Diffusion-7B-Video2World-Sample-MultiView, is fine-tuned from the Diffusion-7B-Video2World-Sample-MultiView model to support video-based conditioning. It achieves this by incorporating previous frames into the generation process. Cosmos-1.0-Diffusion-7B-Video2World-Sample-MultiView can take the video output from Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView and generate its extension. All three models output 6 views of 57 frames of video at a resolution of .
@@
View-dependent cross-attention. In our multi-view setting, each of the six views of the same scene would have a different video description. While we treat all six views as a whole as the state of the diffusion process and perform self-attention among all the elements in the six views for denoising, we find it beneficial to employ view-dependent cross-attention for textual inputs. Specifically, the cross-attention operation for each view only attends to the textual description for the specific view. Note that each view has a different video description in our dataset. With the view embedding and view-dependent cross-attention, we derive Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView from fine-tuning Cosmos-1.0-Diffusion-7B-Text2World.
Trajectory control condition. Optionally, in addition to the text condition, we fine-tune the model to produce videos that conform to the given future trajectory paths to enable more precise control of the agent. This enables the generation of unique driving scenarios that adhere to both driving trajectories recorded by real-world data and driving environments specified by the input text descriptions. The fine-tuned model is Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond.
We define a trajectory as a sequence of 64 points in the 3D space, representing a sequence of translations of the agent from the initial position to the final destination, with each point separated by a 0.1-second interval. We compute the embedding of the trajectory input and make the result a conditional input to the denoiser of the fine-tuned Cosmos-1.0-Diffusion-7B-Video2World model. We note that it is possible to achieve more fine-grained control signals by giving a per-interval action vector, following prior works (Kim et al., 2020, 2021; Hu et al., 2023) or as in the robotic manipulation task (Sec. 6.2). We leave such extensions for future work.

### v2

@@ 5.2.6 Results
In Fig. 17, we show qualitative results of our autoregressive WFMs using different model sizes. In the unprompted setting, comparing Cosmos-Predict1-4B and Cosmos-Predict1-12B model, we observe that the 12B model generates videos with better motion and sharper details. Similarly, in the prompted setting, comparing Cosmos-Predict1-5B-Video2World and Cosmos-Predict1-13B-Video2World reveals that the 13B model gets better motion than the 5B model.
@@
| Output of Cosmos-Predict1-4B | |||
@@
| Cosmos-Predict1-4B | ||
| Cosmos-Predict1-5B-Video2World | ||
| Cosmos-Predict1-12B | ||
| Cosmos-Predict1-13B-Video2World |
@@ 5.2.7 Limitations
One notable failure case observed in the generated videos of our autoregressive WFMs is objects unexpectedly appearing from below. Fig. 19 illustrates an example of this issue. To understand the failure rate of our models, we conduct a systematic study by creating an evaluation set of Physical AI inputs to our autoregressive WFMs. We generate videos with all our models using two input modes—image (single-frame) conditioning and video (9-frame) conditioning. For all generated videos, we manually inspect the failure cases and report the failure rate in Tab. 18. We observe that the smaller models Cosmos-Predict1-4B and Cosmos-Predict1-5B-Video2World show a higher corruption rate in single frame conditioning, while the larger models Cosmos-Predict1-12B and Cosmos-Predict1-13B-Video2World are more robust. Generation with -frame video conditioning is stable for all models, with a failure rate lower than .
@@ 5.3 Evaluation
| Cosmos-Predict1-7B-Text2World | 0.355 | 62.6% | 33.02 | 0.939 | 0.070 |
| Cosmos-Predict1-7B-Video2World | 0.473 | 68.4% | 30.66 | 0.929 | 0.085 |
| Cosmos-Predict1-4B | 0.433 | 35.6% | 32.56 | 0.933 | 0.090 |
| Cosmos-Predict1-5B-Video2World | 0.392 | 27.0% | 32.18 | 0.931 | 0.090 |
@@ 5.3.2 Physics Alignment
| Cosmos-Predict1-7B-Video2World | prompt + 1 frame | 17.34 | 0.538 | 0.836 | 0.332 |
| Cosmos-Predict1-7B-Video2World | prompt + 9 frames | 21.06 | 0.691 | 0.859 | 0.592 |
| Cosmos-Predict1-14B-Video2World | prompt + 1 frame | 16.81 | 0.521 | 0.836 | 0.338 |
| Cosmos-Predict1-14B-Video2World | prompt + 9 frames | 20.21 | 0.635 | 0.860 | 0.598 |
| Cosmos-Predict1-4B | 1 frame | 17.91 | 0.486 | 0.827 | 0.394 |
| Cosmos-Predict1-4B | 9 frames | 18.13 | 0.482 | 0.859 | 0.481 |
| Cosmos-Predict1-5B-Video2World | prompt + 1 frame | 17.67 | 0.478 | 0.818 | 0.376 |
| Cosmos-Predict1-5B-Video2World | prompt + 9 frames | 18.29 | 0.481 | 0.864 | 0.481 |
| Cosmos-Predict1-12B | 1 frame | 17.94 | 0.486 | 0.829 | 0.395 |
| Cosmos-Predict1-12B | 9 frames | 18.22 | 0.487 | 0.869 | 0.487 |
| Cosmos-Predict1-13B-Video2World | prompt + 1 frame | 18.00 | 0.486 | 0.830 | 0.397 |
| Cosmos-Predict1-13B-Video2World | prompt + 9 frames | 18.26 | 0.482 | 0.865 | 0.482 |
@@ 6 Post-trained World Foundation Model
| Sec. 6.1 | Cosmos-Predict1-7B-Video2World-Sample-CameraCond | Text + Image + Cameras |
| Sec. 6.2 | Cosmos-Predict1-7B-Video2World-Sample-Instruction | Text + Video |
| Sec. 6.2 | Cosmos-Predict1-7B-Video2World-Sample-Instruction | Text + Video |
| Sec. 6.2 | Cosmos-Predict1-7B-Video2World-Sample-ActionCond | Action + Video |
| Sec. 6.2 | Cosmos-Predict1-7B-Video2World-Sample-ActionCond | Action + Video |
| Sec. 6.3 | Cosmos-Predict1-7B-Text2World-Sample-MultiView | Text |
| Sec. 6.3 | Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond | Text + Trajectory |
| Sec. 6.3 | Cosmos-Predict1-7B-Video2World-Sample-MultiView | Text + Video |
@@ 6.1 Post-training WFM for Camera Control
Through camera pose conditioning, we integrate camera control into Cosmos-Predict1-7B-Video2World, making it an effective 3D world simulator. We term the result post-trained WFM as Cosmos-Predict1-7B-Video2World-Sample-CameraCond. We focus on generating 3D worlds from a single reference input image, leveraging camera control to produce temporally coherent and 3D-consistent video simulations from the specified camera trajectories, where changes in perspective align with the underlying 3D structure of the scene.
@@ 6.1.2 Fine-tuning
where is the camera center location and is the unit ray direction of each latent pixel (where the latent embedding is treated as a downsampled image). All the camera poses are relative with respect to the initial frame. The Cosmos-Tokenize1-CV888-720p used by Cosmos-Predict1-7B-Video2World models has a temporal compression rate of , and thus for every 8 frames, we use the Plücker embedding at the 4th frame to concatenate with the corresponding latent representation.
@@ 6.1.3 Evaluation
| Cosmos-Predict1-7B-Video2World- | 82.0% | 1.646 | 0.038 | 14.30 | 120.49 |
@@
Comparisons. We present the results in Tab. 22. First, our post-trained WFM can generate realistic and coherent 3D worlds. This is evidenced by the lower FID/FVD scores (higher visual quality) and the higher camera pose estimation success rate. Cosmos-Predict1-7B-Video2World-Sample-CameraCond demonstrates better camera control, as the camera trajectory re-estimation is significantly closer to the original control input.
We also provide visual comparisons in Fig. 21. While CamCo struggles to generate content beyond the input image, Cosmos-Predict1-7B-Video2World-Sample-CameraCond effectively generates visuals that adhere to the structure of a 3D world. Note that both models were post-trained on DL3DV-10K and evaluated on the RealEstate10K dataset, which introduces a significant distribution shift between training and testing. The Cosmos model successfully overcomes this distribution shift while also demonstrating its capability to generalize to unseen input camera trajectories.
@@
To show the diversity of the generation, we show generation results from the same input image and camera control with different random seeds in Fig. 23. Cosmos-Predict1-7B-Video2World-Sample-CameraCond is able to generate different worlds while still maintaining 3D spatial and temporal coherence in the videos. This could be used to simulate different possible futures given the current states.
@@ 6.2.2 Fine-tuning
We fine-tune both our Cosmos-Predict1-7B-Video2World (Sec. 5.1) and Cosmos-Predict1-5B-Video2World (Sec. 5.2) for instruction-based video prediction and action-based next-frame prediction tasks.
For instruction-based video prediction, we build two models based on the base WFMs. The first is called Cosmos-Predict1-7B-Video2World-Sample-Instruction, and the second is called Cosmos-Predict1-5B-Video2World-Sample-Instruction. We compute the T5 embedding of the instruction, which is added to the finetuing of the base model via cross-attention.
For action-based next-frame prediction, we also build two models based on the base WFMs. The first one is called Cosmos-Predict1-7B-Video2World-Sample-ActionCond, and the second one is called Cosmos-Predict1-5B-Video2World-Sample-ActionCond.
Since action is a new modality not encountered during pre-training, we introduce additional modules inside our models for conditioning. For Cosmos-Predict1-5B-Video2World-Sample-ActionCond, we add an action embedder MLP to project the action vector into a tensor, which is then incorporated into the model via cross-attention. For Cosmos-Predict1-7B-Video2World-Sample-ActionCond, we also add an action embedder MLP to predict the action into a tensor but instead, incorporate it into the model by adding it to the timestamp embedding of the DiT modules.
@@ 6.2.3 Evaluation
| Cosmos-Predict1-7B-Video2World-Sample-Instruction | Cosmos-Predict1-5B-Video2World-Sample-Instruction | |||||||||
@@
| Cosmos-Predict1-7B-Video2World-Sample-ActionCond | Cosmos-Predict1-5B-Video2World-Sample-ActionCond | ||||||||||
@@
As shown, we find that both Cosmos-Predict1-7B-Video2World-Sample-Instruction and Cosmos-Predict1-5B-Video2World-Sample-Instruction perform better than VideoLDM-Instruction along the four evaluation dimensions. Cosmos-Predict1-7B-Video2World-Sample-Instruction achieved overall preference compared to for VideoLDM-Instruction. Cosmos-Predict1-5B-Video2World-Sample-Instruction has also achieved better performance than diffusion-based VideoLDM-Instruction. Some predicted video frames for both fine-tuned WFMs are presented in Fig. 25, which shows the quality of the predicted videos.
@@
| Cosmos-Predict1-5B-Video2World- Sample-ActionCond | 19.95 | 0.80 | 0.36 | 434 |
| Cosmos-Predict1-7B-Video2World- Sample-ActionCond | 21.14 | 0.82 | 0.32 | 190 |
The computed metrics are summarized in Tab. 23, including PSNR, SSIM, Latent L2 (Zhu et al., 2024), and FVD. As shown, both Cosmos-Predict1-5B-Video2World-Sample-ActionCond and Cosmos-Predict1-7B-Video2World-Sample-ActionCond models outperform the baseline model (IRASim-Action). Some predicted video frames are presented in Fig. 26, which shows the quality of the predicted videos compared to the ground truth.
@@ 6.3.2 Fine-tuning
We fine-tune our Cosmos-Predict1-7B-Text2World (Sec. 5.1) into a multiple-view world model using the RDS dataset. To ensure consistent video generation across multiple views, we slightly modify the architectural design described in Sec. 5.1 and fine-tune the WFM to generate videos from all six cameras simultaneously.
We build three multi-view world models, summarized in Tab. 21. The first one is called Cosmos-Predict1-7B-Text2World-Sample-MultiView, which is a multi-view world model that can generate six camera views based on a text prompt input. The second one is called Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond. This model is built on top of Cosmos-Predict1-7B-Text2World-Sample-AV-MultiView and takes an additional trajectory input as the conditional input signal. The final model, Cosmos-Predict1-7B-Video2World-Sample-MultiView, is fine-tuned from the Diffusion-7B-Video2World-Sample-MultiView model to support video-based conditioning. It achieves this by incorporating previous frames into the generation process. Cosmos-Predict1-7B-Video2World-Sample-MultiView can take the video output from Cosmos-Predict1-7B-Text2World-Sample-MultiView and generate its extension. All three models output 6 views of 57 frames of video at a resolution of .
@@
View-dependent cross-attention. In our multi-view setting, each of the six views of the same scene would have a different video description. While we treat all six views as a whole as the state of the diffusion process and perform self-attention among all the elements in the six views for denoising, we find it beneficial to employ view-dependent cross-attention for textual inputs. Specifically, the cross-attention operation for each view only attends to the textual description for the specific view. Note that each view has a different video description in our dataset. With the view embedding and view-dependent cross-attention, we derive Cosmos-Predict1-7B-Text2World-Sample-MultiView from fine-tuning Cosmos-Predict1-7B-Text2World.
Trajectory control condition. Optionally, in addition to the text condition, we fine-tune the model to produce videos that conform to the given future trajectory paths to enable more precise control of the agent. This enables the generation of unique driving scenarios that adhere to both driving trajectories recorded by real-world data and driving environments specified by the input text descriptions. The fine-tuned model is Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond.
We define a trajectory as a sequence of 64 points in the 3D space, representing a sequence of translations of the agent from the initial position to the final destination, with each point separated by a 0.1-second interval. We compute the embedding of the trajectory input and make the result a conditional input to the denoiser of the fine-tuned Cosmos-Predict1-7B-Video2World model. We note that it is possible to achieve more fine-grained control signals by giving a per-interval action vector, following prior works (Kim et al., 2020, 2021; Hu et al., 2023) or as in the robotic manipulation task (Sec. 6.2). We leave such extensions for future work.